OpenStudy (anonymous):

Solve for x: 3(x-2) = - 4(x+8) + 10

OpenStudy (watchmath):

\(3x-6=-4x-32+10\) \(3x+4x=6+10-32\) \(7x=-16\) \(x=-16/7\)

OpenStudy (anonymous):

3x - 6 = -4x - 32 +10 3x + 4x = 10-32+6 7x = -16 Therefore x = -16/7

OpenStudy (anonymous):

i didn't have -4 thats why my answes was incorrect

OpenStudy (anonymous):

thanks!